On 21 January 2019 banang invested an amount in an account earning 6.5% simple interest. If she has R1200 available on 9 May 2019 then the amount that she invested on 21 Jan was


1
Expert's answer
2021-11-23T17:31:44-0500

Simple interest=principal×R×T


Interest=1200


Rate=6.5


Time=3.6/12


P=??


1200=p×6.5/100×3.6/12


Make p the subject


P=1200/0.0195


P=61538.462